Tests made on the breaking strength of 10 pieces of a metal gave the following results: 578, 572, 570, 568, 572, 570, 570, 572, 598 and 584 kg. Test if the mean breaking strength of the wire can be assumed as 577 kg


1
Expert's answer
2022-04-11T10:58:57-0400

"H_0: \mu=577" kg

"H_a: \mu\neq577" kg


"t=\frac{\overline{x}-\mu}{s\sqrt{n}}"


"\overline{x}=\sum{x_i}/n=575.4" kg


"s=\sqrt{\frac{\sum (x_i-\overline{x})|^2}{n-1}}=9.24" kg


"t=\frac{575.4-577}{9.24/\sqrt{10}}=-3.162"


"df=n-1=10-1=9"


critical value for "\alpha=0.05" :

"t_{crit}=2.262"


Since "|t|>t_{crit}" , we reject the null hypothesis.

The mean breaking strength of the wire cannot be assumed as 577 kg